# Covariance spectrum of MAXI J1820+070: On the nature of the Comptonizing flow

### v1

@@ 1 Introduction
In systems with accreting black holes (BHs), such as X-ray binaries (XRBs) that contain stellar-mass black holes and active galactic nuclei (AGNs) housing supermassive black holes, the observed X-ray photons primarily originate from the regions near the black holes. In the so-called hard state, the X-ray energy spectrum typically features two components: weak thermal emission in the low-energy band and strong power-law emission in the high-energy band (Zdziarski2004). The weak thermal emission is believed to originate from a geometrically thin, optically thick disk, while the strong power-law emission is attributed to inverse Comptonization occurring close to the black hole (Galeev1979; Haardt1991; Svensson1994).
1.1 Models of the accretion flow
@@
It has been proposed that the thin disk is truncated at a specific radius, within which a hot flow develops (Esin1997). The co-evolution of the thin disk and the hot flow plays a crucial role in the state transitions observed during the outbursts of XRBs. Recently, you2023 demonstrated that, in the framework of a truncated disk model, the truncation radius of MAXI J1820+070 increases as the mass accretion rate declines, indicating that the inner hot flow is expanding as this source transitions from the soft state to the hard state. In this context, soft photons originating from the outer disk can act as seed photons. These photons are Compton up-scattered by thermal electrons within the hot flow through inverse Compton processes. This mechanism accounts for the observed power-law-like X-ray continuum.
@@
The nature of the Comptonizing hot flow is still a topic of debate. On the one hand, the advection-dominated accretion flow (ADAF), proposed by Narayan1994, has been widely used to explain the hard X-ray emission observed in BHXRBs, often referred to as a corona. This corona is commonly regarded as the primary region for Comptonization. On the other hand, past studies have suggested that the jet base could also be a potential site for Comptonization emission (Markoff2005). In this context, either the disk blackbody or synchrotron photons could serve as the dominant seed photons for the inverse-Compton scattering process. In MAXI J1820+070, during the rising hard state, i.e., before transitioning to the soft state, the illumination of the hard X-ray Comptonization emission on the disk is weakening according to the broad X-ray spectral fits (you2021), as the hard X-ray Comptonization is contracting towards the BH (kara2019; DeMarco2021). It was hypothesized that the source of X-ray Comptonization may act as an outflowing corona during the rising hard state of MAXI J1820+070, contributing to the observed decrease in its illumination (you2021).
When simultaneous observations in optical and/or infrared wavelengths are coupled with radio and X-ray data, a broadband spectral energy distribution (SED) can be constructed. The analysis of these fits will enhance our understanding of the sources of X-ray Comptonization (marino2021; rodi2021; ozbey2022; banerjee2024; et2024; yoshitake2022; Tantry2025). It has been observed that during the hard state, jet emission is characterized by a flat to slightly inverted, optically thick spectrum that extends from the radio to the infrared/optical bands (Markoff2001; Corbel2002; carotenuto2021a). The jet spectrum transitions from optically thick to optically thin emission with a spectral break at $\nu_{b}\sim 10^{11–14}\rm{Hz}$ (Russell2013a; Russell2013b). This transition may also significantly contribute to the infrared/optical bands (Markoff2003; russell2006; Tetarenko2015; Du2025). The optical and infrared (OIR) photons can serve as seed photons for Comptonization.
1.2 Measurements and nature of variability
@@
In addition to the multi-wavelength observation on a long timescale of days, as mentioned above, the fast variability on a timescale of seconds in the inter-bands is also crucial in revealing the physics of the X-ray emission (Belloni2014; Bollemeijer2025; Bollemeijer2025b; gao2025; Xu2025; Zhan2025) and the accretion–ejection connection (Tetarenko2021; Paice2021; Vincentelli2021; Zhang2022; Bu2021). One can construct X-ray variability spectra by calculating the PDS for each energy channel and integrating it over a specific frequency range to obtain the variance in that channel. This yields the root-mean-square (RMS) spectrum, which characterizes the variability of spectral components within that frequency range (Vaughan2003). Wilkinson2009 introduced the technique of covariance spectra, which shares similarities with RMS spectra but is more robust in low signal-to-noise regimes. It measures the covariance between a given energy channel and a broader reference band, enabling the isolation of variability contributions from different spectral components on distinct timescales. Covariance spectra and ratios derived from XMM-Newton observations of GX 339-4 revealed enhanced disk blackbody variability relative to the Comptonized emission below one keV on timescales longer than one second, which are caused by instabilities in the disc itself. As a result, we aim to extend the covariance spectrum into the high-energy band, enabling us to study X-ray Comptonization from a timing perspective. It is essential to note that the power-law Comptonization spectrum decreases at higher energies, with a cutoff that reaches approximately 100 keV. This cutoff is typical for the hard spectral states observed in BHXRBs (You2023ApJ). Therefore, it is crucial to ensure that X-ray observations extend into the high-energy band.
1.3 Detailing properties of MAXI J1820+070
@@
In 2018, the outburst of MAXI J1820+070 was discovered in X-rays by the Monitor of All-sky X-ray Image (MAXI) on March 11. The long-term and high-frequency observation of MAXI J1820++070 by the Insight-HXMT took place from March to October 2018. During this period, Insight-HXMT monitored the outburst over 140 times, accumulating a total exposure of 2560 ks. The high statistics for high-energy photons and the broad energy range (1–250 keV) of Insight-HXMT enable detailed timing analysis of high-energy and broad-band variability. In this study, we focus on the variability properties of broadband noise (BBN). We utilize data from Insight-HXMT to investigate the broadband X-ray variability of the black hole X-ray binary MAXI J1820+070 during its hard state. By applying covariance spectral analysis to the variability over different timescales, we explore the behavior and evolution of BBN across a wide energy range. Our objective is to probe the possible presence of multiple Comptonization regions. More details on the data used and on the data analysis are given in Sect. 2. In Sect. 3, we present the broadband coherence and covariance ratios on different timescales, and characterize the observed variability through modeling of the covariance spectra. In Sect. 4, to interpret the observed variability properties, we explore the possible origins of the hard X-ray emission.
2 Observations and Data Reduction
@@
The X-ray outburst of MAXI J1820+070 began with a rapid increase in flux, peaking around MJD 58200. This was followed by a gradual decline that lasted until approximately MJD 58290. Subsequently, the source underwent a re-brightening between MJD 58290 and MJD 58305. The hard X-ray flux then decreased dramatically around MJD 58305, indicating a transition from a hard to a soft state. After remaining in the soft state for more than two months, from MJD 58305 to MJD 58380, the hard X-ray flux began to increase around MJD 58380 as the source transitioned back to the hard state. Figure 1 illustrates the Insight-HXMT light curves for this outburst, observed by the low-energy (LE, 1 to 10 keV), medium-energy (ME, 10 to 30 keV), and high-energy (HE, 27 to 150 keV) instruments. In this study, we focus on the spectral-timing analysis of hard X-ray emissions from the corona, utilizing over 140 Insight-HXMT observations that capture the source during its initial hard state (MJD 58191-58300). Due to the enormity of the dataset, in this paper, we select five representative epochs from the total Insight-HXMT observations, as shown in Figure 1, to illustrate our results. The information of the selected epochs is detailed in Table 1.
@@
3 Data Analysis and Results
3.1 X-ray timing analysis
@@
Figure 2: Power density spectrum (PDS) of each epoch in the hard state. The color scheme of the observations is consistent with that shown in Figure 1. Successive spectra are vertically offset by 0.03 for clarity. The PDS are derived from the 2–10 keV light curves obtained with the LE detector of Insight-HXMT. The shaded regions denote the frequency ranges corresponding to the long- and short-timescale broadband noise in each epoch.
3.1.1 Power density spectra
@@
We utilize the software Stingray for timing analysis of X-ray data (Huppenkothen2019) to calculate the PDS. The lightcurve is divided into continuous intervals of 100 seconds, with a time resolution of 0.01 seconds, which corresponds to a Nyquist frequency $f_{\text{Nyq}}=50$ Hz. Then, we compute periodograms for each interval and average them across observations. The PDS is normalized using rms-squared normalization, specifically to the fractional RMS (Miyamoto1991). These PDS reveal significant variability across a wide frequency range observed by Insight-HXMT, extending from soft to hard X-rays.
@@
The Poisson statistics of the power spectrum, denoted as $P_{\text{noise}}$, is calculated using the formula $P_{\text{noise}}=\langle\Delta x^{2}\rangle/(\langle x\rangle^{2})f_{\text{Nyq}}$, where $\langle\Delta x^{2}\rangle$ represents the average of the squared error bars of the light curve. For the observations made by Insight-HXMT, the dead time $\tau_{d}$ is approximately 20 µs for both HE and LE bands, and 250 µs for the ME band. The frequency range typically analyzed in stellar-mass BH systems is well below $1/\tau_{d}$. Given the minimal dead time for HE and LE, our timing results will not be affected by the deadtime. For the ME band, the deficit in the power density due to deadtime is approximately 1.9%, which is less than the background uncertainty of about 3%. Therefore, the effect of dead time can be reliably ignored (Ma2021).
@@
During the hard state from MJD 58191 to MJD 58300, a low-frequency QPO was detected in 76 observations of Insight-HXMT (Ma2021). The QPO centroid frequency $f$ varied between 0.03 and 0.7 Hz. The quality factor $Q$, defined as the ratio of the QPO centroid frequency $f$ to its full width at half maximum $\Delta f$, was determined to be greater than 2.5 (Ma2021). To investigate the variability of the BBN, we excluded variability from both the QPO and its harmonic. In each observation that featured the QPO centroid frequency $f$, we selected two frequency ranges: 0.010.7f0.01-0.7f Hz, referred to as long timescale noise, and 2.3f5.0f2.3f-5.0f Hz, referred to as short timescale noise. These frequency bands correspond to the flat top and the declining portion of the PDS, respectively. Between MJD 58192 and MJD 58197, no QPO signals were detected. To maintain consistency with the other data, we applied the same long- and short-timescale frequency ranges to the data collected during this stage as used for the observation on MJD 58197 (obsid: P0114661005). An example of the resulting PDS, after subtracting Poisson noise, is shown in Figure 2.
@@
Figure 3: Coherence as a function of energy over the long timescale (left panel) and short timescale (right panel) frequency ranges for our five observation epochs. The coherence spectra are derived from the five epochs of Insight-HXMT data, including the rise and decay of the hard state outburst. The reference band is selected as 2–10 keV.
3.1.2 Energy-dependent coherence
@@
The derived PDS shows that both the magnitude and shape of the noise component change significantly with energy (Yang2022). To further examine the variability of the broadband noise, we first measure the intrinsic coherence $\gamma^{2}$ as a function of energy (Vaughan_1997; Uttley2014), using the 2-10 keV band as the reference band. For each light curve across the 50 energy channels, we calculate the intrinsic coherence between it and the reference light curve on both long and short timescales, following the procedure outlined in Uttley2014.
@@
The coherence function quantifies the degree of linear correlation between two light curves within a specified frequency range (Vaughan_1997; Uttley2014). The intrinsic coherence as a function of energy, derived on both long and short timescales for the selected epochs, is shown in Figure 3. It is found that the coherence changes with energy. On long timescales, as shown in the left panel of Figure 3, the coherence remains stable around unity for energies between 2 and 10 keV, and begins to decline with energy in high-energy bands (>30>30keV). This indicates that hard X-ray variability at energies above $\sim 30$ keV is incoherent with soft X-ray variability in the 2-10 keV range. On short timescales, a similar trend is exhibited, as illustrated in the right panel of Figure 3. Note that the incoherence is also observed in other epochs of Insight-HXMT data throughout the hard state on both long and short timescales. The evolution of intrinsic coherence with energy suggests at least two distinct variability components in the hard X-ray band. One component varies linearly with the soft band (2-10 keV) component, dominating at tens of keV, while the other varies incoherently with the soft component, dominating at higher energies on both timescales.
3.1.3 The root mean square and covariance spectra

### v2

@@ I Introduction
In systems with accreting black holes (BHs), such as X-ray binaries (XRBs) that contain stellar-mass black holes and active galactic nuclei (AGNs) housing supermassive black holes, the observed X-ray photons primarily originate from the regions near the black holes(Remillard and McClintock, 2006; Done et al., 2007). In the so-called hard state, the X-ray energy spectrum typically features two components: weak thermal emission in the low-energy band and strong power-law emission in the high-energy band (Zdziarski and Gierliński, 2004; Belloni, 2010; Done et al., 2007). The weak thermal emission is believed to originate from a geometrically thin, optically thick disk(Shakura and Sunyaev, 1973). The observed hard X-ray power-law continuum is commonly attributed to Comptonization of soft seed photons by hot electrons in the inner accretion flow(Sunyaev and Titarchuk, 1980; Paczynski, 1978; Svensson and Zdziarski, 1994; Melia and Misra, 1993; Done et al., 2007).
I.1 Models of the accretion flow
@@
The advection-dominated accretion flow (ADAF), proposed by Narayan and Yi (1994, 1995), has been widely used to explain the hard X-ray emission observed in BHXRBs, often referred to as a corona. This corona is commonly regarded as the primary region for Comptonization. The hot flow are postulated to exist within a specific radius, where the outer thin disk is truncated (Esin et al., 1997; Done et al., 2007). The co-evolution of the thin disk and the hot flow plays a crucial role in the state transitions observed during the outbursts of XRBs(Esin et al., 1997; Belloni, 2010; Done et al., 2007). Esin et al. (1997) proposed that the inner thin disk radius increases during the soft-to-hard state transitions. Recently, You et al. (2023a) demonstrated that, in the framework of a truncated disk model, the truncation radius of MAXI J1820+070 increases as the mass accretion rate declines, indicating that the inner hot flow is expanding as this source transitions from the soft state to the hard state. In this context, soft photons originating from the outer disk can act as seed photons. These photons are Compton up-scattered by thermal electrons within the hot flow through inverse Compton processes. This mechanism accounts for the observed power-law X-ray continuum.
@@
However, the nature of the Comptonizing hot flow is still a topic of debate. Alternative models have suggested that synchrotron emission may also account for the broad continuum features observed in XRBs (Markoff et al., 2001, 2005b, 2005a). And the synchrotron photons could also serve as the seed photons for the inverse-Compton scattering process (Fabian et al., 1982; Markoff et al., 2005b; Veledina et al., 2011, 2013). In MAXI J1820+070, during the rising hard state, i.e., before transitioning to the soft state, the illumination of the hard X-ray Comptonization emission on the disk is weakening according to the broad X-ray spectral fits (You et al., 2021), as the hard X-ray Comptonization is contracting towards the BH (Kara et al., 2019; De Marco et al., 2021). It was hypothesized that the source of X-ray Comptonization may act as an outflowing corona during the rising hard state of MAXI J1820+070, contributing to the observed decrease in its illumination (You et al., 2021). Discriminating among different physical scenarios for the hard state requires strong observational constraints. Both spectral and timing properties should be considered(Uttley et al., 2014).
I.2 Measurements and nature of variability
@@
In addition to the multi-wavelength observation on a long timescale of days, as mentioned above, the fast variability on a timescale of seconds in the inter-bands is also crucial in revealing the physics of the X-ray emission (van der Klis, 2006; Uttley et al., 2014; Belloni and Stella, 2014; Bollemeijer et al., 2025; Bollemeijer and Uttley, 2025; Gao et al., 2025; Xu et al., 2025; Zhan et al., 2025) and the accretion–ejection connection (Tetarenko et al., 2021; Paice et al., 2021; Vincentelli et al., 2021; Zhang et al., 2022; Bu et al., 2021). One can construct X-ray variability spectra by calculating the PDS for each energy channel and integrating it over a specific frequency range to obtain the variance in that channel. This yields the root-mean-square (RMS) spectrum, which characterizes the variability of spectral components within that frequency range (van der Klis, 2006; Vaughan et al., 2003). Wilkinson and Uttley (2009) introduced the technique of covariance spectra, which shares similarities with RMS spectra but is more robust in low signal-to-noise regimes. It measures the covariance between a given energy channel and a broader reference band, enabling the isolation of variability contributions from different spectral components on distinct timescales. Covariance spectra and ratios derived from XMM-Newton observations of GX 339-4 revealed enhanced disk blackbody variability relative to the Comptonized emission below one keV on timescales longer than one second(Wilkinson and Uttley, 2009; Uttley et al., 2011), which are caused by instabilities in the disc itself. As a result, we aim to extend the covariance spectrum into the high-energy band, enabling us to study X-ray Comptonization from a timing perspective. It is essential to note that the power-law Comptonization spectrum decreases at higher energies, with a cutoff that reaches approximately 100 keV. This cutoff is typical for the hard spectral states observed in BHXRBs (Gierliński et al., 1997; Zdziarski and Gierliński, 2004; You et al., 2023b). Therefore, it is crucial to ensure that X-ray observations extend into the high-energy band.
I.3 Detailing properties of MAXI J1820+070
@@
In 2018, the outburst of MAXI J1820+070 was discovered in X-rays by the Monitor of All-sky X-ray Image (MAXI) on March 11(Kawamuro et al., 2018). The long-term and high-frequency observation of MAXI J1820++070 by the Insight-HXMT took place from March to October 2018(You et al., 2021). During this period, Insight-HXMT monitored the outburst over 140 times, accumulating a total exposure of 2560 ks. The high statistics for high-energy photons and the broad energy range (1–250 keV) of Insight-HXMT enable detailed timing analysis of high-energy and broad-band variability(Kara et al., 2019; Buisson et al., 2019; De Marco et al., 2021). In this study, we focus on the variability properties of broadband noise (BBN). We utilize data from Insight-HXMT to investigate the broadband X-ray variability of the black hole X-ray binary MAXI J1820+070 during its hard state(Zdziarski et al., 2021; Stiele and Kong, 2020). By applying covariance spectral analysis to the variability over different timescales, we explore the behavior and evolution of BBN across a wide energy range. Our objective is to probe the possible presence of multiple Comptonization regions(Ingram and Done, 2012; Veledina et al., 2013; Zdziarski et al., 2021). More details on the data used and on the data analysis are given in Sect. II. In Sect. III, we present the broadband coherence and covariance ratios on different timescales, and characterize the observed variability through modeling of the covariance spectra. In Sect. IV, to interpret the observed variability properties, we explore the possible origins of the hard X-ray emission.
II Observations and Data Reduction
@@
The X-ray outburst of MAXI J1820+070 began with a rapid increase in flux, peaking around MJD 58200. This was followed by a gradual decline that lasted until approximately MJD 58290. Subsequently, the source underwent a re-brightening between MJD 58290 and MJD 58305. The hard X-ray flux then decreased dramatically around MJD 58305, indicating a transition from a hard to a soft state. After remaining in the soft state for more than two months, from MJD 58305 to MJD 58380, the hard X-ray flux began to increase around MJD 58380 as the source transitioned back to the hard state. Figure 1 illustrates the Insight-HXMT light curves for this outburst, observed by the low-energy (LE, 1 to 10 keV), medium-energy (ME, 10 to 30 keV), and high-energy (HE, 27 to 150 keV) instruments. In this study, we focus on the spectral-timing analysis of hard X-ray emissions from the corona, utilizing over 140 Insight-HXMT observations that cover the source during its initial hard state (MJD 58191–58300). Due to the enormity of the dataset, in this paper, we select five representative epochs from the total Insight-HXMT observations, as shown in Figure 1, to illustrate our results. Before MJD 58230, the HE count rates in the 28-150 keV are above $2\times 10^{3}$ cnt/s, providing high-quality timing results. Therefore, all five selected epochs are drawn from this period. Observations after MJD 58230 show qualitatively similar behavior, but with poorer statistical quality due to the lower count rates. The information of the selected epochs is detailed in Table 1.
@@
III Data Analysis and Results
III.1 X-ray timing analysis
@@
Figure 2: Power density spectrum (PDS) of each epoch in the hard state. The color scheme of the observations is consistent with that shown in Figure 1. The PDS are derived from the 2–10 keV light curves obtained with the LE detector of Insight-HXMT. The centroid QPO frequencies for the five epochs are 0.029, 0.042, 0.067, 0.081, 0.130 Hz, respectively. The shaded regions denote the frequency ranges corresponding to the long- and short-timescale broadband noise in each epoch.
III.1.1 Power density spectra
@@
We utilize the software Stingray for timing analysis of X-ray data (Huppenkothen et al., 2019) to calculate the PDS. The lightcurve is divided into continuous intervals of 100 seconds, with a time resolution of 0.01 seconds, which corresponds to a Nyquist frequency $f_{\text{Nyq}}=50$ Hz. Then, we compute periodograms for each interval and average them across observations. The PDS is normalized using rms-squared normalization, specifically to the fractional RMS (Miyamoto et al., 1991). These PDS reveal significant variability across a wide frequency range observed by Insight-HXMT, extending from soft to hard X-rays.
@@
The Poisson statistics of the power spectrum, denoted as $P_{\text{noise}}$, is calculated using the formula $P_{\text{noise}}=\langle\Delta x^{2}\rangle/(\langle x\rangle^{2})f_{\text{Nyq}}$, where the variable x represents the photon count rate. $\langle\Delta x^{2}\rangle$ represents the average of the squared error bars of the light curve. For the observations made by Insight-HXMT, the dead time $\tau_{d}$ is approximately 20 µs for both HE and LE bands, and 250 µs for the ME band. The frequency range typically analyzed in stellar-mass BH systems is well below $1/\tau_{d}$. Given the minimal dead time for HE and LE, our timing results will not be affected by the deadtime. For the ME band, the deficit in the power density due to deadtime is approximately 1.9%, which is less than the background uncertainty of about 3%. Therefore, the effect of dead time can be reliably ignored (Ma et al., 2021). The resulting PDS of the five selected epochs, after subtracting Poisson noise, are shown in Figure 2.
@@
During the hard state from MJD 58191 to MJD 58300, a low-frequency quasi-periodic oscillation (QPO) was detected in 76 observations of Insight-HXMT (Ma et al., 2021). The QPO centroid frequency $f$ varied between 0.03 and 0.7 Hz. The quality factor $Q$, defined as the ratio of the QPO centroid frequency $f$ to its full width at half maximum $\Delta f$, was determined to be greater than 2.5 (Ma et al., 2021). In this work, our primary goal is to investigate the BBN properties of the source. For this reason, we first excluded variability associated with both the QPO and its harmonic. For each observation in which a QPO with centroid frequency $f$ is detected, we selected two frequency ranges: 0.010.7f0.01-0.7f Hz, referred to as long timescale, and 2.3f5.0f2.3f-5.0f Hz, referred to as short timescale. Based on the measured FWHM of the QPO, these two frequency bands are expected to be dominated by BBN. They correspond to the flat-top and declining portions of the PDS, respectively. Between MJD 58192 and MJD 58197, no QPO signals were detected. To maintain consistency with the other data, we applied the same long- and short-timescale frequency ranges to the data collected during this stage as used for the observation on MJD 58197 (obsid: P0114661005).
@@
Figure 3: Coherence as a function of energy over the long timescale (left panel) and short timescale (right panel) frequency ranges for our five observation epochs. The coherence spectra are derived from the five epochs of Insight-HXMT data, including the rise and decay of the hard state outburst. The 2–10 keV band is selected as the reference.
III.1.2 Energy-dependent coherence
@@
The derived PDS across soft and hard X-ray energies show that both the magnitude and shape of the noise component change significantly with energy (Yang et al., 2022). To further examine the energy dependence of the BBN variability, we first measure the intrinsic coherence $\gamma^{2}$ as a function of energy (Vaughan and Nowak, 1997; Uttley et al., 2014), using the 2-10 keV band as the reference band. For each light curve across the 50 energy channels, we calculate the intrinsic coherence between it and the reference light curve on both long and short timescales, following the procedure outlined in Uttley et al. (2014). For further details, please refer to the Appendix.
@@
The coherence function quantifies the degree of linear correlation between two light curves within a specified frequency range (Vaughan and Nowak, 1997; Uttley et al., 2014). The intrinsic coherence as a function of energy, derived on both long and short timescales for the selected epochs, is shown in Figure 3. We find that the coherence changes with energy. On long timescales, as shown in the left panel of Figure 3, the coherence remains stable around unity in the 2-10 keV range, as expected since this band is used as the reference. In the 10–30 keV range, the variability also remains coherent, indicating that the soft X-ray emission varies in a similar linear manner across these energies. Interestingly, the coherence begins to decline with energy in hard X-ray bands (>30>30keV). This behavior suggests that the hard X-ray variability above $\sim 30$ keV becomes increasingly incoherent with respect to the soft X-ray variability in the 2–10 keV band. On short timescales, a similar trend is exhibited, as illustrated in the right panel of Figure 3. Note that the incoherence is also observed in other epochs of Insight-HXMT data throughout the hard state on both long and short timescales. The evolution of intrinsic coherence with energy suggests the presence of at least two distinct variability components in the X-ray band. One component varies linearly with the soft band (2-10 keV) component, dominating at tens of keV, while the other varies incoherently with the soft component, dominating at higher energies on both timescales.
III.1.3 The root mean square and covariance spectra